OpenStudy (anonymous):

The sum of two numbers is 66. The larger number is three less than twice the smaller number. Find the numbers.

OpenStudy (anonymous):

x = Small Number y = Large Number By the question, y = 2x-3 So: x+y=66 Substitute Y for 2x-3... \[x+(2x-3)=66\] and solve for x: \[3x - 3 = 66\] \[3x=69\] \[x = 23\] 23 is your small number, now plus that back in to the large number question (y). \[y=2(23)-3=43\] Small Number: 23 Large Number: 43 Check if needed: 43+23=66